Circular No. 5969

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 R. Chini, H.-P. Gemuend, G. Haslam, E. Kreysa, P. Lambertz, and

R. Zylka, Max-Planck-Institut für Radioastronomie, Bonn; and R. Tuffs, Max-Planck-Institut für Kernphysik, Heidelberg, report the detection of SN 1994I at 1.3 mm using the IRAM 30-m telescope (+ MPIfR 7-element bolometer array): “Observations indicate that the emission at this wavelength is essentially constant over 2 days. The central beam of the hexagonal array was pointed towards the optical position of SN 1994I (IAUC 5961). No discrete sources were detected in the six surrounding beams, which are offset 22′′ from the central beam. The angular resolution was 11′′ (FWHM). The following flux densities were derived by subtracting the average of flux in the surrounding beams from the central one; they therefore represent the excess flux at the position of the supernova over the diffuse emission from M51: Apr. 6.07 UT, 9.0 ± 2.0 mJy; 7.00, 8.4 ± 0.6; 8.04, 8.2 ± 1.0. The errors denote 1σ random uncertainties. The calibration of the absolute flux scale, which was made using Uranus (assuming a temperature of 101 K for the planet’s disk), introduces an additional systematic uncertainty of 10 percent. The 1.3-mm measurements lie below an extrapolation in epoch and observing frequency of the VLA measurements (IAUC 5966). Optical-depth effects, inverse Compton processes near maximum light, and intrinsic particle acceleration time scales in the source could all contribute to this different behavior. Monitoring at 1.3 mm will continue.”

PU VULPECULAE P. Garnavich, Dominion Astrophysical Observatory, writes: “A spec-

trogram (range 465–505 nm) of PU Vul was obtained with the DAO 1.2-m telescope on Apr. 1.5 UT. The broad He II 468.6-nm emission feature, seen for the past four years, is absent. Compared with a spectrum obtained in July 1993, the strength of He II is down by at least a factor of ten relative to Hβ and supports the suggestion that an eclipse of the hot source is in progress (IAUC 5960). The equivalent width of Hβ has increased between these two observations, which implies that the continuum around 480 nm has become weaker. The nebular lines of [O III] at 495.9/500.7 nm have strengthened relative to Hβ by ∼ 15 percent since July 1993, but this may be unrelated to the eclipse and due to the slowly increasing temperature of the source.”

1994 April 8 Daniel W. E. Green